# Discovering Properties of Inflectional Morphology in Neural Emergent Communication

### v1

@@ Abstract
Emergent communication (EmCom) with deep neural network-based agents promises to yield insights into the nature of human language, but remains focused primarily on a few subfield-specific goals and metrics that prioritize communication schemes which represent attributes with unique characters one-to-one and compose them syntactically. We thus reinterpret a common EmCom setting, the attribute-value reconstruction game, by imposing a small-vocabulary constraint to simulate double articulation, and formulating a novel setting analogous to naturalistic inflectional morphology (enabling meaningful comparison to natural language communication schemes). We develop new metrics and explore variations of this game motivated by real properties of inflectional morphology: concatenativity and fusionality. Through our experiments, we discover that simulated phonological constraints encourage concatenative morphology, and emergent languages replicate the tendency of natural languages to fuse grammatical attributes.
Discovering Properties of Inflectional Morphology in Neural Emergent Communication
@@ 1 Introduction
Emergent communication (EmCom) with deep neural agents studies communication protocols developed between two or more agents engaged in a task requiring successful communication. Analyses of learned protocols in EmCom typically search for characteristics deemed crucial in natural language, prominently compositionality Chaabouni et al. (2020). Such experiments promise to shed light on the pressures underlying human-like language evolution.
@@
In this work, we reinterpret EmCom as emergent morphology, addressing these shortcomings with two key design choices. First, we target humanlike double articulation Hockett (1960), the property of natural language that meaningful components are composed of combinations of individually meaningless components, by enforcing a small character inventory for messages, thus necessitating the combination of characters into meaningful morpheme-like symbols to express attribute values. Second, as shown in Figure 1, we introduce a novel configuration of attribute-value reconstruction inspired by inflectional morphology, where a high-cardinality attribute (analogous to roots) is paired with comparatively lower-cardinality attributes (analogous to grammatical information, e.g., tense and person), encouraging richer morphological phenomena. We review EmCom evaluation metrics for various properties of communication schemes, proposing novel metrics for the underexplored properties of concatenativity and fusionality. To validate and establish expected ranges for these metrics, we then analyze a representative group of artificial double articulatory communication schemes and inflection-based natural language communication schemes.
Lastly, we leverage this reimagined problem to explore several novel, linguistically compelling research questions. We implement a phonology-inspired constraint in communication agents, using concatenativity metrics to reveal ease of articulation as a pressure toward concatenative languages. We then compare the topographic similarity of emergent and natural languages, finding significant room for improvement in emergent language to reflect natural degrees of compositionality. Finally, we analyze the presence of fusion in emergent and natural languages for inflection, finding that emergent languages most often fuse grammatical attributes rather than roots, mirroring natural language. Altogether, this work lays a strong foundation for future research into the evolution of morphology.
@@ 2 Related Work
We adopt the term “emergent morphology” introduced by Archangeli and Pulleyblank (2016) within their Emergent Grammar framework, which hypothesizes how morphophonological structures arise from constraint interactions in natural languages. This is one of few attempts to model the emergence of aspects of morphological systems, including work by Elsner et al. (2020), who simulate how phonological processes can lead agglutinative systems to become fusional over time. Also related are formal and evolutionary analyses of double articulation conducted by Nowak and Krakauer (1999) and Zuidema and de Boer (2009). Unlike these works, we emphasize computational modeling of artificial communication systems, focusing specifically on how minimal meaningless discrete units combine systematically into meaningful morphological structures like inflection.
There exist some works that address double articulation explicitly via selection of small inventory sizes Ueda et al. (2023); Ueda and Washio (2021); Ueda and Taniguchi (2024), but the analysis focuses on words rather than morphemes, and only considers perfectly concatenative methods of combining morphemes. Work utilizing bitstring messages in EmCom also achieve this Resnick et al. (2020); Gupta et al. (2020), albeit not in a linguistically realistic way.
Other works use large-vocabulary EmCom to investigate various phenomena relevant to natural language morphology. For example, Chaabouni et al. (2019) use a length pressure to rederive Zipf’s Law of Abbreviation, a statistical property of natural language. Lian et al. (2023) investigate the specific morphosyntactic problem of the case-marking versus word-order tradeoff with pre-learned languages as a starting point, with Lian et al. (2024) extending this inquiry to a group communication setting. Conklin and Smith (2023) discover naturalistic variability in emergent languages, complicating traditional measures of compositionality. While Galke et al. (2022) argue the importance of learnability, generalization, and group size pressures in EmCom, Galke and Raviv (2024) further review the impact of learnability and production effort among other psycho-sociolinguistic factors in EmCom. Meanwhile, Dekker (2024) models the evolution of specific languages’ inflectional morphology by pretraining agents with those languages, whereas we focus on the emergence of inflectional morphology systems from scratch.
Lastly, there is extensive literature on automated morphological segmentation, and segmenters for nonconcatenative morphology are particularly relevant Fullwood (2018). While we only explored concatenative approaches, as emergent morphology develops, we expect these techniques to play a crucial role in understanding emergent languages.
@@ 3 Toward Emergent Morphology
Prior work in EmCom has largely focused on the Lewis Signaling Game Kottur et al. (2017); Chaabouni et al. (2020); Ueda et al. (2023); Lewis (1979), in which a sender agent observes some state and communicates via a message to a receiver agent, which must take an action based on this information. Success in the game relies on accurate communication of the relevant features of the environment. Morphology has been relatively understudied in EmCom due to a focus on individual symbols representing atomic units of meaning Boldt and Mortensen (2024b).
@@ 3.1 Attribute-Value Reconstruction Game
Formally, an instance of Attr-Val can be defined as such: we have a state , a sender , and a listener . is the message space, which consists of messages of up to length characters, all of which must belong to a fixed character vocabulary . This instance is considered successful if . Following prior works, we use EGG Kharitonov et al. (2019) to implement sender and receiver agents as single layer gated recurrent unit (GRU; Cho et al., 2014) models. More details about agent implementation and training are provided in Appendix F.
3.2 Attr-Val Configurations as Inflectional Morphology
The attribute value sets tend to have the same cardinality in prior work Kottur et al. (2017); Chaabouni et al. (2020); Ueda et al. (2023). So, as a conceptual framework, this setting has commonly been compared to communicating about objects with various properties e.g., a red square or blue triangle.
We claim that Attr-Val games also represent a rich analogy for inflectional morphology. The meaning of an inflected root in natural language consists of the root as well as additional attributes like tense and person for verbs, or number and gender for nouns. Like Attr-Val, these attributes take on a finite set of values. While the number of possible roots is unbounded in natural language, Attr-Val can simulate this by having one attribute set (representing roots) be much larger than others. This paradigm provides a key advantage: natural languages solve an analogous and similarly challenging problem to this, enabling direct comparison with learned communication schemes.
We consider two configurations for , both with three attribute sets: a default setting with 16 possible values per attribute (comparable to prior work), and a novel inflection setting with attribute set cardinalities of 42, 2, and 3, respectively symbolizing roots, tenses, and person.111Note that we also conducted preliminary experiments controlling for total number of combinations (683 2 3), but models universally failed to converge, leading us to conclude that symbolic complexity Zhang (2025) is a more meaningful and necessary quantity to control for. We further discuss this issue and justify our choices of values in Appendix I. Both settings have a comparable total number of attribute values (48 and 47). To necessitate double articulation, i.e., the composition of meaningful symbols from meaningless characters to represent attributes, we enforce a consistent character vocabulary size and message length , smaller than the cardinality of most attribute sets. Importantly, one attribute in the inflection setting has a cardinality much greater than while others have fewer possible values than the vocabulary size, possibly incentivizing learning of more morphologically diverse communication schemes than the default setting.
@@ 4 Interpreting Communication Schemes
To investigate the emergence of double articulation in EmCom, we next introduce methods to evaluate communication schemes by extracting meaningful symbols from messages, and measuring various properties of them. Toward establishing a typology of possible compositional communication schemes, we then establish reference points for emergent communication schemes based on both idealized artificial languages and natural languages.
@@ 4.1.2 Evaluation Metrics
Given a segmented symbol vocabulary , we next introduce common metrics for its compositionality from prior work, along with methods to evaluate concatenativity and fusionality, more specific forms of compositionality that typical compositionality metrics do not directly measure.
@@
However, TopSim places minimal restrictions on how similarity of messages is calculated. As such, Chaabouni et al. (2020) propose bag-of-symbols disentanglement (BoSDis), an alternative metric accounting for permutation-invariant communication schemes222They also propose positional disentanglement, which we omit but provide more details about in Appendix B. (e.g., in English, “red and blue” is equivalent to “blue and red”):
BoSDis calculates the average difference in mutual information between the count of a vocabulary symbol and attributes and , which respectively have the most and second-most mutual information with . This is normalized by entropy for . This metric captures the idea that a symbol’s presence in a message uniquely communicates a specific value of a specific attribute. This should generally not be true in our small-vocabulary setting, which necessitates reuse of characters in multi-character symbols to induce double articulation,333An obvious exception is one-character symbols, which are possible but can only comprise a small proportion of symbols in an accurate communication scheme. thus BoSDis should generally be low. However, if a symbol segmentation algorithm is effective at identifying symbols which uniquely refer to some attribute, the BoSDis with respect to those symbols should be higher. We thus use this as a gauge of whether HAS and BPE extract meaningful symbols through a ratio BoSDis of the BoSDis for a segmented symbol vocabulary to that for the unsegmented character vocabulary . If BoSDis, this indicates that creates a more disentangled representation than individual characters , suggesting successful identification of meaningful multi-character symbols.
@@
TopSim and BoSDis fail to account for non-trivial and variable forms of compositionality, including some present in natural language Korbak et al. (2020); Conklin and Smith (2023). For example, TopSim based on Levenshtein distance of characters reflects an assumption that characters should not be modified or reordered under composition. Contrarily, we aim to consider all schemes in which messages are formed via a systematic composition of symbols (i.e., as opposed to a memorized “hash function”) as compositional.
@@
One particular form of compositionality we aim to tease apart from others is concatenativity, i.e., the degree to which composable symbols consist of uninterrupted streams of characters (analogous to phonemes) that are concatenated under composition. 444This property is closely related to the constraint of integrity in optimality theory, which prohibits output forms from containing multiple correspondents to an input form McCarthy and Prince (1995). Using English verb conjugation as an example, “played” is a concatenative message consisting of “play” and “ed.” By contrast, “sang” is a nonconcatenative message since the characters “s_ng” correspond to the root “sing,” but are interrupted by “a,” which communicates tense.
@@
As HAS and BPE require symbols to be composed of consecutive characters, we estimate concatenativity as the average number of symbols in segmented messages of a communication scheme.555For HAS, we use the number of boundaries drawn. Note that concatenative compositionality implies compressibility via algorithms like BPE, but not necessarily vice versa. Since HAS relies on entropy, symbol length based on it may appear artificially low in communication schemes that include some random or otherwise non-meaningful characters. We refer to these length metrics as HASLen and BPELen|V|. A smaller number of symbols indicates greater concatenativity, as more consecutive characters belong in self-contained units of meaning. Of course, not all compositional communication schemes are concatenative. In our forthcoming experiments, we will apply concatenativity metrics to contrived concatenative and nonconcatenative compositional schemes (introduced in Section 5.1) to analyze how these properties relate.
Fusionality
@@
Fusionality or degree of fusion is the expression of multiple attributes of meaning within one morpheme or symbol. Prior work examines this by taking conditional probabilities of surface forms after ablating feature pairs from the dataset Rathi et al. (2021, 2022). However, this requires training a model for each feature pair, a computationally expensive approach. The fusion of two attributes is equivalent to treating them as one attribute , for which symbols may be chosen arbitrarily, and a language may be compositional with respect to a particular fused attribute set. As such, we define a novel metric of fused TopSim (F-TopSim), calculated by the maximum TopSim over the fusions of all attribute pairs. A language is fusional if , which we summarize by a F-TopSim .
@@ 5 Artificial and Natural Reference Points
Next, we curate reference compositional communication schemes. To evaluate the ability of metrics to capture compositionality, concatenativity, and fusionality and establish signatures as expected ranges of metric, we first propose several contrived artificial languages to target these properties. To contextualize EmCom results with real-world communication schemes, we develop additional languages based on natural language inflections.
@@ 5.1 Artificial Languages
As shown in Figure 2, we generate five artificial languages using identical message length and character vocabulary size to Attr-Val settings. The first four types compose symbols with various levels of concatenativity, considering symbols of constant and mixed length. Notably, these languages preserve symbols and the ordering of their characters under composition, favored by TopSim. The fusion language, however, violates this by assigning symbols to combinations of attribute values rather than individual ones. These languages illustrate a range of morphological complexity.666Additional artificial languages explored in Appendix C. In Table 1, we report metrics for the compositionality (including effectiveness of symbol extraction), fusionality, and learnability of artificial languages. In Table 2, we report concatenativity metrics. We provide an extended discussion and additional artificial languages in Appendix C, summarizing key observations below.
@@
First, the degree of concatenation in artificial languages has minimal effect on learnability, implying a lack of inductive bias toward concatenation in recurrent networks. BoSDis metrics indicate that HAS and BPE extract more meaningful morphemes from more concatenative languages, despite all artificial languages being compositional. HASLen and BPELen are lower in more concatenative languages, suggesting our proposed metrics effectively reflect concatenativity. Mixed concatenative languages have higher variance in BPELen, falling along a spectrum between perfect concatenation and nonconcatenation, and random languages have markedly higher asymptotic BPELen than systematic languages, further evidence of BPE’s effectiveness. However, as expected, HAS is less effective for languages with randomness, evidenced by mixed concatenative and variable length languages having relatively low HASLen and HAS BoSDis, and random languages having low HASLen.777Supplementary results with HASLen in Appendix D. BoSDis metrics themselves are consistently greater than 1 for random languages, suggesting an over-sensitivity to noise in the absence of other signal. Lastly, F-TopSim effectively reflects the fusionality of the fusion language.

### v2

@@ Abstract
Emergent communication (EmCom) with deep neural network-based agents promises to yield insights into the nature of human language, but remains focused primarily on a few subfield-specific goals and metrics that prioritize communication schemes which represent attributes with unique characters one-to-one and compose them syntactically. We thus reinterpret a common EmCom setting, the attribute-value reconstruction game, by imposing a small-vocabulary constraint to simulate double articulation, and formulating a novel setting analogous to naturalistic inflectional morphology (enabling meaningful comparison to natural language communication schemes). We develop new metrics and explore variations of this game motivated by real properties of inflectional morphology: concatenativity and fusion. Through our experiments, we discover that simulated phonological constraints encourage concatenative morphology, and emergent languages replicate the tendency of natural languages to fuse grammatical attributes.
Discovering Properties of Inflectional Morphology
in Neural Emergent Communication
@@ 1 Introduction
Emergent communication (EmCom) studies communication protocols developed between two or more deep neural agents engaged in a task requiring successful communication. Analyses of learned protocols in EmCom typically search for characteristics deemed crucial in natural language, prominently compositionality Chaabouni et al. (2020). Such experiments promise to shed light on the pressures underlying human-like language evolution.
@@
In this work, we reinterpret EmCom as emergent morphology, addressing these shortcomings with two key design choices. First, we target humanlike double articulation Hockett (1960) by enforcing a small character inventory, necessitating that individually meaningless characters form combinatorial morpheme-like symbols to encode attribute values and combinations thereof. Second, as shown in Figure 1, we introduce a novel configuration of attribute-value reconstruction inspired by inflectional morphology, where a high-cardinality attribute (analogous to roots) is paired with comparatively lower-cardinality attributes (analogous to grammatical information, e.g., tense and person), encouraging richer morphological phenomena. We review EmCom evaluation metrics for various properties of communication schemes, proposing novel metrics for the underexplored properties of concatenativity and degree of fusion. To validate and establish expected ranges for these metrics, we then analyze a representative group of artificial double articulatory communication schemes and inflection-based natural language communication schemes.
Lastly, we leverage this reimagined problem to explore novel, linguistically compelling research questions. We implement a phonology-inspired constraint in communication agents, using concatenativity metrics to reveal ease of articulation as a pressure toward concatenative languages. We then compare the topographic similarity of emergent and natural languages, finding significant room for improvement in emergent language to reflect natural degrees of compositionality. Finally, we analyze the presence of fusion in emergent and natural languages for inflection, finding that emergent languages most often fuse grammatical attributes rather than roots, mirroring natural language. This work opens exciting new directions for future research into the evolution of morphology.
@@ 2 Related Work
Some other EmCom works address double articulation explicitly via selection of small inventory sizes Ueda and Washio (2021); Ueda et al. (2023); Ueda and Taniguchi (2024), but the analysis focuses on words rather than morphemes, and only considers perfectly concatenative methods of combining them. Work utilizing bitstring messages in EmCom achieves this Resnick et al. (2020); Gupta et al. (2020), but not in a linguistically realistic way.
Unlike our work, many works use large-vocabulary EmCom to investigate various phenomena relevant to natural language morphology. For example, Chaabouni et al. (2019) use a length pressure to rederive Zipf’s Law of Abbreviation. Lian et al. (2023, 2024) explore the case-marking versus word-order tradeoff in pair and group communication. Conklin and Smith (2023) discover naturalistic variability in emergent languages through proposed metrics for synonymy, homonymy, word-order freedom, and entanglement, complicating typical measures of compositionality. Galke et al. (2022); Galke and Raviv (2024) explore the importance of learnability, generalization, production effort, and group size pressures, among other psycho-sociolinguistic factors. Chaabouni et al. (2022) further explore the relationships between these aspects in a naturalistic visual task setting, while Kouwenhoven et al. (2024) investigate the effect of agents’ representational alignment on their performance.
Lastly, there is vast literature studying morphology in human languages. Notably, we adopt the term “emergent morphology” introduced by Archangeli and Pulleyblank (2016) within their Emergent Grammar framework, which hypothesizes how morphophonological structures arise from constraint interactions in natural languages. Especially related prior works attempt to model and analyze the emergence of various aspects of morphological systems Nowak and Krakauer (1999); Zuidema and de Boer (2009); Elsner et al. (2020); Dekker (2024). Unlike these works, we emphasize computational modeling of artificial communication systems, focusing on how minimal meaningless discrete units combine systematically into meaningful morphological structures like inflection. Automated morphological segmentation, especially for nonconcatenative morphology Fullwood (2018), is also relevant. While we only explored concatenative approaches, we expect such techniques to play a crucial role in future work toward understanding morphology in EmCom.
@@ 3 Toward Emergent Morphology
Prior EmCom work has largely focused on the Lewis Signaling Game Kottur et al. (2017); Chaabouni et al. (2020); Ueda et al. (2023); Lewis (1979), in which a sender agent observes some state and generates a message to a receiver agent, which must take an action based on it. Success in the game relies on accurately communicating the relevant features of the environment. Morphology has been relatively understudied in EmCom due to a focus on individual characters representing atomic units of meaning Boldt and Mortensen (2024b).
@@ 3.1 Attribute-Value Reconstruction Game
Formally, Attr-Val can be defined as such: we have a state , a sender , and a listener . is the message space, which consists of messages of up to length characters, all of which must belong to a fixed character vocabulary . This instance is considered successful if . Following prior works, we use EGG Kharitonov et al. (2019) to implement sender and receiver agents as single layer gated recurrent unit (GRU; Cho et al., 2014) models.111More details about agents provided in Appendix A.
3.2 Attr-Val as Inflectional Morphology
The attribute value sets tend to have the same cardinality in prior work Kottur et al. (2017); Chaabouni et al. (2020); Ueda et al. (2023), often motivating analogies to communicating properties of objects with few possible values, e.g., a red square or blue triangle. We claim that Attr-Val games also represent a rich analogy for inflectional morphology. The meaning of an inflected root in natural language consists of the root as well as additional attributes like tense and person for verbs, or number and gender for nouns. Like Attr-Val, these attributes take on a finite set of values. While the number of possible roots is unbounded in natural language, Attr-Val can simulate this through one attribute set (representing roots) being much larger than others. This paradigm provides a key advantage: natural languages solve an analogous and similarly challenging problem to this, enabling direct comparison with learned communication schemes.
We consider 4 configurations for : a default setting resembling prior work with 3 attributes of 16 values each , and 3 novel inflection settings with nonuniform attribute value sets , , and . In inflection settings, the largest attribute set symbolizes roots, while others symbolize grammatical properties like tense and person. Settings have a comparable total number of attribute values (47-48).222We also ran experiments controlling for total number of combinations, but models failed to converge (see Appendix B). To necessitate double articulation, i.e., composing meaningless characters into meaningful symbols to represent attributes, we enforce a consistent character vocabulary size and message length , smaller than the size of most attribute sets. Importantly, one attribute in the inflection settings has many more possible values than while others have fewer, possibly incentivizing more morphologically diverse communication schemes.
@@ 4 Interpreting Communication Schemes
To investigate the emergence of double articulation in EmCom, we next introduce methods to evaluate communication schemes by extracting meaningful symbols from messages, and measuring various properties of them. We then establish reference points for emergent communication schemes based on both idealized artificial languages and natural languages.
@@ 4.1.2 Evaluation Metrics
Given a segmented symbol vocabulary , we next discuss common metrics for its compositionality from prior work. We then introduce methods to evaluate concatenativity and degree of fusion, more specific forms of compositionality that typical compositionality metrics do not directly measure, as well as learnability, which has previously been connected to compositionality Kirby et al. (2014); Li and Bowling (2019); Chaabouni et al. (2020).
@@
However, TopSim places minimal restrictions on how similarity of messages is calculated. As such, Chaabouni et al. (2020) propose bag-of-symbols disentanglement (BoSDis), an alternative metric accounting for permutation-invariant communication schemes333They also propose positional disentanglement, which we omit but provide more details about in Appendix D. (e.g., “A and B” versus “B and A”):
BoSDis calculates the mean entropy-normalized difference in mutual information between the count of a vocabulary symbol and attributes and , which have the 2 highest values with . This measures whether a character’s presence in a message uniquely communicates a specific attribute value. This should generally not happen in our small-vocabulary setting, which necessitates reuse of characters in multi-character symbols to induce double articulation,444One-character symbols are an exception, but an accurate communication scheme has very limited capacity for these. thus BoSDis should be low. However, if an algorithm effectively segments symbols which uniquely refer to attributes, the BoSDis with respect to those symbols should be higher. We thus gauge whether HAS and BPE extract meaningful symbols through a ratio BoSDis of BoSDis for a segmented symbol vocabulary and unsegmented characters . If BoSDis, then creates a more disentangled representation than individual characters , thus contains meaningful multi-character symbols.
@@
TopSim and BoSDis fail to account for non-trivial and variable forms of compositionality, including some present in natural language Korbak et al. (2020); Conklin and Smith (2023). For example, TopSim based on Levenshtein distance of characters reflects an assumption that characters should not be modified or reordered under composition. Contrarily, we aim to consider any scheme where messages are formed via a systematic composition of symbols (i.e., as opposed to a memorized “hash function”) as compositional.
@@
One particular form of compositionality we aim to tease apart from others is concatenativity, i.e., the degree to which composable symbols consist of uninterrupted streams of characters (analogous to phonemes) that are concatenated under composition.555This is closely related to integrity in optimality theory, which prohibits output forms from containing multiple correspondents to an input form McCarthy and Prince (1995). Using English verb conjugation as an example, “played” is a concatenative message consisting of “play” and “ed.” By contrast, “sang” is a nonconcatenative message since the characters “s_ng” correspond to the root “sing,” but are interrupted by “a,” which communicates tense.
@@
As HAS and BPE require symbols to be composed of consecutive characters, we estimate concatenativity as the average number of symbols in segmented messages of a communication scheme. We refer to these length metrics as HASLen and BPELen|V|.666For HASLen, we use the number of boundaries. Since HAS relies on entropy, we note that HASLen may be artificially low in communication schemes that include some random or otherwise non-meaningful characters. A smaller number of symbols indicates greater concatenativity, as more consecutive characters belong in self-contained units of meaning. Of course, not all compositional communication schemes are concatenative. In Section 5.1, we will apply concatenativity metrics to contrived concatenative and nonconcatenative compositional schemes to analyze how these properties relate.
Degree of Fusion
@@
Fusion is the expression of multiple attributes of meaning within one morpheme or symbol. Prior work examines this by taking conditional probabilities of surface forms after ablating feature pairs from data Rathi et al. (2021, 2022), but this requires computationally expensive training of models for each feature pair. The fusion of two attributes is equivalent to treating them as one attribute , for which symbols may be chosen arbitrarily, and a language may be compositional with respect to a particular fused attribute set. As such, we define a novel metric of fused TopSim (F-TopSim), calculated by the maximum TopSim over the fusions of all attribute pairs. A language is fusional if , so we summarize the degree of fusion with F-TopSim .
Learnability
To explore its relationship with concatenativity and fusion, we measure learnability by the average number of training epochs that a listener takes to exceed 99% reconstruction accuracy.
@@ 5 Artificial and Natural Reference Points
Next, we curate reference compositional communication schemes. To evaluate how metrics capture the above properties and establish expected ranges, we first propose artificial languages to target them. To contextualize EmCom results with real-world communication schemes, we develop additional languages based on natural language inflections.
@@ 5.1 Artificial Languages
As shown in Figure 2, we generate 5 artificial languages with a range of morphological complexity using message length and vocabulary size . The first 4 types compose symbols of constant and mixed length with varied levels of concatenativity. These languages preserve symbols and the ordering of their characters under composition, favored by TopSim. The fusion language violates this by assigning symbols to combinations of attribute values rather than individual ones. Tables 1-2 report metrics for listeners trained on these languages. We list key observations below.777More discussion and artificial languages in Appendix E.
First, the degree of concatenation has minimal effect on learnability, implying a lack of inductive bias toward concatenation in recurrent listener networks. BoSDis metrics confirm that HAS and BPE extract more meaningful morphemes from more concatenative languages, despite all languages being compositional. HASLen and BPELen are lower in more concatenative languages, suggesting these metrics effectively reflect concatenativity. Mixed concatenative languages have higher variance in BPELen, falling along a spectrum between perfectly concatenative and nonconcatenative languages, and random languages have markedly higher asymptotic BPELen than others, further evidence of BPE’s effectiveness. As expected, HAS is less effective for languages with randomness, evidenced by low HAS BoSDis and HASLen in mixed concatenative and variable length languages, with the latter also low in random languages.888Supplementary results with HASLen in Appendix F. BoSDis metrics are large for random languages, suggesting an over-sensitivity to noise in the absence of other signal. Lastly, F-TopSim effectively reflects fusion in the fusion languages.